Potter Clarkson
All Systems Operational

About This Site

Welcome to the Potter Clarkson status page, all of our systems are constantly monitored and issues reported here. Scheduled maintenance times, and the areas affected, are also displayed.

Use the 'subscribe to updates' button at the top of this page to add your work or personal email address and receive notifications of problems and planned maintenance schedules.

R3 Operational
K2 Operational
iManage Operational
Aderant Expert Operational
EPO Online Filing System Operational
Office 365 Operational
Mimecast Operational
VMWare View Operational
VPN and Network Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Apr 1, 2020

No incidents reported today.

Mar 31, 2020
Completed - The scheduled maintenance has been completed.
Mar 31, 19:30 BST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Mar 31, 19:01 BST
Update - We will be undergoing scheduled maintenance during this time.
Mar 31, 08:44 BST
Scheduled - This evening we will be applying the latest EPO software update to the EPO Online Filing System. During this time you will not be able to use the application.
Mar 31, 08:43 BST
Mar 30, 2020

No incidents reported.

Mar 29, 2020

No incidents reported.

Mar 28, 2020

No incidents reported.

Mar 27, 2020

No incidents reported.

Mar 26, 2020

No incidents reported.

Mar 25, 2020
Completed - The scheduled maintenance has been completed.
Mar 25, 20:00 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Mar 25, 19:02 GMT
Scheduled - Virtual computers will be refreshed this evening to bring them back to normal performance levels.
Mar 25, 18:44 GMT
Resolved - This incident has been resolved.
Mar 25, 18:47 GMT
Update - We are continuing to work on a fix for this issue.
Mar 25, 18:45 GMT
Update - We are continuing to work on a fix for this issue.
Mar 25, 18:44 GMT
Update - Good news. The faulty server is now online and working as normal. At 7pm this evening all virtual computers will be restarted to bring performance back to standard operating levels.
Mar 25, 17:36 GMT
Update - Replacement parts have arrived and have now been installed. We are working to bring the server back online as quickly as possible.
Mar 25, 14:42 GMT
Update - We are currently awaiting delivery of replacement parts for the failed hardware. This is scheduled from between 10am and 2pm today. Once the parts are here and working we'll update this status message.
Mar 25, 10:29 GMT
Identified - We are currently running on a single VMWare server due to a disk failure, parts should be with us today. Status updates will appear on this page.
Mar 25, 09:00 GMT
Mar 24, 2020
Resolved - This incident has been resolved.
Mar 24, 17:39 GMT
Identified - A server is currently off line and the team are investigating.
Mar 24, 14:07 GMT